Corporation A has $84973 in taxable income, and Corporation B has $9.3 million in taxable income. Suppose both firms have identified a new project that will increase taxable income by $11550. How much more will Corporation B pay in additional taxes than will Corporation A? Use the tax rates in Table 2.3 TABLE 2.3 Taxable Income Tax Rate Corporate Tax Rates 50,000 75,000 100,000 100,001335,000 335,001-10,000,000 10,000,001-15,000,000 15,000,001-18,333,333 15% 25 34 39 34 35 38 35 50,001 75,001 18,333,334+
Answer #1

Firstly, we write down the tax slab rate table in standard form -

If taxable income is over.. But not over… The Tax is … Marginal Tax rate
$                            -   $                        50,000 15% of the amount over $ 0 15%
$                  50,001 $                        75,000 $7,500 plus 25% of the amount over $ 50,000 25%
$                  75,001 $                      100,000 $13,750 plus 34% of the amount over $ 75,000 34%
$                100,001 $                      335,000 $22,250 plus 39% of the amount over $ 100,000 39%
$                335,001 $                10,000,000 $113,900 plus 34% of the amount over $ 335,000 34%
$          10,000,001 $                15,000,000 $3,400,000 plus 35% of the amount over $ 10,000,000 35%
$          15,000,001 $                18,333,333 $ 5,150,000 plus 38% of the amount over $ 15,000,000 38%
$          18,333,334 $6,416,666.54 Plus 35% of the amount over $ 18,333,334 35%

with this we will calculate the tax liability of both companies before new project.

Corporation A - Tax liability before new project

A's Taxable income is $ 84,973

Thus, A's Tax liability = $ 13,750 + 0.34(84,973-75,000)

= $ 13,750 + $ 3,390.82

= $ 17,140.82

Corporation B - Tax liability before new project

B's Taxable Income = $ 9,300,000

Thus, B's Tax Liability = $ 113,900 + 0.34 ( 9,300,000 - 335,000)

= $ 3,162,000

Corporation-A - Tax liability after new project

A's taxable income = 84,973 + 11,550 = $ 96,523

Thus, A's Tax liability = 13,750 + 0.34(96,523 -75,000)

= $ 21,067.82

A's Additional tax for new project = 21,067.82 - 17,140.82 = $ 3,927

Corporation-B - Tax liability after new project

B's Taxable income = 9,300,000 + 11,550 = $ 9,311,550

B's Tax liability = 113,900 + 0.34 (9,311,550 - 335,000)

= $ 3,165,927

B's Additional tax for new project = 3,165,927 - 3,162,000 = $ 3,927

Additional tax liability of Both Corporations A and B due to new project is same i.e 3,927.

Thus, Corporation-B is not paying additional tax in excess of Corporatation-A

This is because both corporation A & B comes in a Tax bracket having equal Marginal tax rate i.e 34%